OpenStudy (anonymous):

Select the choice below that could be the first step in solving the following system of equations using elimination: -6x - 7y = 12 3y - 4x -1 = 0 Answer Replace the right side of the first equation with the second equation. Multiply the second equation by 4. Multiply the second equation by -3. Write the second equation in terms of y. @uri can you help me?

OpenStudy (texaschic101):

-6x - 7y = 12 -4x + 3y = 1 well...you could multiply the 2nd equation by -3 making 12x, and then the 1st equation by 2, making -12x....eliminating the x's out